OpenStudy (anonymous):

chris's hike started at an elevation of -12 m. His hike finished at an elevation of 235 m. How much did chris's elevation change? m

OpenStudy (anonymous):



OpenStudy (anonymous):

ill give medal and fan

OpenStudy (anonymous):

223 meters

OpenStudy (anonymous):

any way thanks

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

@lostintime is it right

OpenStudy (anonymous):

223 m is it right

OpenStudy (anonymous):

I'm not very good at math sorry but here is someone who might be able to help: @robtobey

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@Slhics next time try to explain your answer, giving direct answers is not permitted on this site and you can get in trouble, not trying to be mean but I am trying to caution you from any further problems.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Black_ninja123 @hhelpplzzzz @KendrickLamar2014

OpenStudy (anonymous):

thx any way @lostintime

OpenStudy (black_ninja123):

I think you have to subtract 235 - -12

OpenStudy (anonymous):

Yes @black_ninja 123

OpenStudy (anonymous):

If that's true and you subtract 235-(-12) then I got 247

OpenStudy (anonymous):

i think it cause when you have -(- you turn it into a plus sign and it becomes 235+12 rather than subtracting...

OpenStudy (black_ninja123):

yes 235 - -12 is 247

OpenStudy (anonymous):

awesome I did it right thanks for confirming @Black_ninja123

OpenStudy (anonymous):

so its 247

OpenStudy (black_ninja123):

Yes and Np @lostintime

OpenStudy (anonymous):

nice @lostintime

OpenStudy (anonymous):

Yes I believe so but I'd like to know how @Slhics got 223 meters? hm

OpenStudy (anonymous):

235 - (-12) = 235 + 12 = 247

OpenStudy (black_ninja123):

I think she forgot the -12 and she did positive 12 instead

OpenStudy (anonymous):

Well @robtobey @Black_ninja123 and I all got the same answer so it must be correct :] oh yeah that makes sense

OpenStudy (anonymous):

maybe she thought the (-) was the subtraction sign well either way hope you got the help you needed @WOLFPACK13 good luck!

OpenStudy (anonymous):

thx

OpenStudy (black_ninja123):

If you need help with anything else just tag us

OpenStudy (anonymous):

ok

OpenStudy (mathmate):

or do a vertical number line:  so to go from -12 to 235, you would have to go first 12 m to zero, and then add it to from zero to 235, so total height climbed would be 12+235=247. Using @robtobey's method would be even faster, final height - initial height = 235-(-12)=247

OpenStudy (anonymous):

thats what we did... 235-(-12) = 235 + 12 = 247

OpenStudy (mathmate):

Just to say that the number line is helpful if you want to double check if you should add or subtract.